MS F-stat Among-group Within Total 243.87548.775 280.38 Total number of observations: 63 (a). The number of different groups (c) is: (b). Make the hypothesis statement for this ANOVA test: He i Not all s are equal (c). The mean square for within-group (MSW) is (Three decimals): (d). The F statistic is (Three decimals): (e). The critical value, Fa, at α-0.05 is (Three decimals):
Answer #1

Total no. of observation = 63

a) Number of different groups, c= df Among +1

df_{Among}= SSA/ MSA

= 243.875 / 48.775 = 5

Number of different groups, c= 5+1 = 6

b) Null and Alternative Hypothesis:

H_o:\mu_1=\mu_2=\mu_3=\mu_4= \mu_5=\mu_6

H_1: \textup{Not all mean are equal.}

c) df_{total} = total-1 = 63 -1= 62

df_{within}= df_{total}-df_{among} = 62 - 5 = 57

MSW = SSW/ df_{within} = 280.38/ 57 = 4.919

d) F = MSA / MSW

= 48.775/ 4.919 = 9.916

e) At \alpha = 0.05, critical value F_\alpha = 2.376